Victor has a piece of wood that is 56 cm long. He cuts the wood into two pieces so that the ratio of the legnth of the shorter piece is 3:4. What is the length of the longer piece?
Answer by erika514(39)

Okay.....
So we are going to set up an equation to go to the situation. Our equation is going to be 3x+4x=56. The solution for x will be multiplied by the coefficients to help find the length of the two pieces of wood. So, lets start solving...
First, we are going to combine our x terms to get 7x. We should now have 7x=56 as our equation. Finally to get our solution for x, we will divide each side by 7 to get simply x on the one side and 56 divided by 7 on the other side which is just 8. So your value for x is 8. We can now plug in 8 for x to find the lengths of the pieces of wood. So we need to put 8 in for all our x terms in the equation 3x+4x=56-->3(8)+4(8)=56. So our pieces are 24 and 32 cm. in length. Let's check to make sure we ar correct. So we have 24 and 32 for our lengths, so 24+32 is 56, and 56 equals 56. So we have the right lengths. Finally we look at our two values and determine that 32 cm. is longer than 24 cm., so our longer piece has a length of 32 cm. Good luck, let me know if you have any questions. :)